# Frozen Large Language Models Can Perceive Paralinguistic Aspects of Speech

### v1

@@ IV-C Training and evaluation
TABLE II: Responses of Baseline 2 and SpeechEmotionLlama v3 to the speech prompt “I can’t believe I ripped my pants when I came over the wall though” spoken in the “laughing” style. Note that the baseline does not perceive or acknowledge the laughing nature of the speech in its response, while SpeechEmotionLlama does.
@@
v3 “Oh no, that’s hilarious! I can just imagine the look on your face! Did you have to do a quick fix or did you just laugh it off and keep going? What was going on that made you jump over the wall so suddenly? Was it a dare or a sudden impulse?”
V Results
V-A Responses to expressive speech prompts
Table I shows the performance of the various models in terms of the quality of their responses to expressive speech prompts. Baseline 2 obtained much better scores than Baseline 1, thanks to having been fine-tuned on in-domain data. However, its emotion understanding score was still quite low; this is unsurprising, since its training should not have allowed it to learn or utilize any paralinguistic information from speech inputs. Cascading both baselines with the SER model resulted in marginal gains in terms of response quality and emotion understanding, but the improvements were quite small.
Modifying training to perform alignment against emotion-conditioned responses resulted in large improvements in terms of all metrics. ROUGE, BERTScore, and LLM-annotated quality scores all increased for SpeechEmotionLlama v1, but perhaps the most notable improvements were in emotion understanding, which improved by over 1.7 points compared to the cascaded SER + Baseline 2 system. Adding additional style and emotion-related response alignment tasks (v2 and v3) did not lead to significant changes to either of the LLM-annotated scores, but provided slight improvements to ROUGE and BERTScore. These results indicate that our proposed training framework enabled SpeechEmotionLlama to take into account the emotion or speaking style of a speech prompt and respond accordingly, unlike baselines that were not trained on emotion-conditioned responses. Table II shows examples of responses from Baseline 2 and SpeechEmotionLlama v3 to an expressive speech prompt demonstrating this.
V-B Speech emotion recognition
To more explicitly examine how well each system’s speech tokens convey emotion/style information to the LLM, we prompted the different model variants to perform SER by prepending the same prompt as used in the SER training task before the speech tokens. Then, we checked whether the response generated by the LLM matched the emotion/style label. The rightmost column of Table I shows these results; the only exceptions are the cascade models, for which we list the performance of the standalone SER model. As expected, the two baselines performed quite poorly. Fine-tuning on emotion-conditioned response alignment resulted in a significant increase in performance, and adding speech captioning and SER each resulted in further improvements. Surprisingly, SpeechEmotionLlama v3 even outperformed the standalone SER model, achieving 81.51% accuracy. This further demonstrates that our training framework allows the speech encoder to effectively convey paralinguistic information to the LLM in a way that allows the LLM to make use of that information.
@@ V-B Speech emotion recognition
V-C Analysis of speech tokens
We performed extra analyses to study the amount of paralinguistic information encoded in the different systems’ speech tokens. We first trained linear probes on temporally pooled speech tokens from each model variant to perform the 15-way SER task. The two baselines obtained 25.08% and 31.20% accuracy, while SpeechEmotionLlama v1, v2, and v3 obtained higher accuracies of 68.15%, 73.15%, and 78.40%, respectively. We also performed a t-SNE [34] analysis of the mean-pooled speech tokens. Fig. 2 displays the resulting plot, which shows that the speech tokens from SpeechEmotionLlama are clustered much more clearly by emotion/style tag compared to the baseline. These results demonstrate that training on more emotion-related response alignment tasks produces token embeddings that encode more discriminative paralinguistic information from speech.
VI Conclusion
We study the potential of an LLM to understand paralinguistic aspects of speech without any fine-tuning of its weights. Building on prior work that equips frozen LLMs with the ability to process speech inputs, we propose a novel training paradigm called behavioral alignment against emotion-conditioned responses. Our framework enables a speech encoder to generate token embeddings that capture both semantic and paralinguistic information from a spoken utterance and effectively convey it to the LLM, even when the LLM remains completely frozen. The LLM, in turn, is able to effectively utilize the paralinguistic information in the speech tokens to provide higher quality and more empathetic responses to expressive speech prompts. Our work opens up new possibilities for developing interfaces to LLMs that can leverage their capabilities for other kinds of speech or audio tasks without the need for any fine-tuning.

### v2

@@ 3.3 Training and evaluation
Table 2: Responses of Baseline 2 and SpeechEmotionLlama v3 (“SEL v3”), to the speech prompt “I can’t believe I ripped my pants when I came over the wall though” spoken in the “laughing” style. Note the bold text; the baseline does not perceive or acknowledge the laughing nature of the speech in its response, while SpeechEmotionLlama does.
@@
SEL v3 “Oh no, that’s hilarious! I can just imagine the look on your face! Did you have to do a quick fix or did you just laugh it off and keep going? What was going on that made you jump over the wall so suddenly? Was it a dare or a sudden impulse?”
4 Results
4.1 Responses to expressive speech prompts
Table 1 shows the performance of the various models in terms of the quality of their responses to expressive speech prompts. Baseline 2 obtained much better scores than Baseline 1, thanks to having been fine-tuned on in-domain data. However, its emotion understanding score was still quite low; this is unsurprising, since its training should not have allowed it to learn or utilize any paralinguistic information from speech inputs. Cascading both baselines with the SER model resulted in marginal gains in terms of response quality and emotion understanding, but the improvements were quite small.
Modifying training to perform alignment against emotion-conditioned responses resulted in large improvements in terms of all metrics. ROUGE, BERTScore, and LLM-annotated quality scores all increased for SpeechEmotionLlama v1, but perhaps the most notable improvements were in emotion understanding, which improved by over 1.7 points compared to the cascaded SER + Baseline 2 system. Adding additional style and emotion-related response alignment tasks (SpeechEmotionLlama v2 and v3) did not lead to significant changes to either of the LLM-annotated scores, but provided slight improvements for ROUGE and BERTScore. These results indicate that our proposed training framework enabled SpeechEmotionLlama to take into account the emotion or speaking style of a speech prompt and respond accordingly, unlike baselines that were not trained on emotion-conditioned responses. Table 2 shows examples of responses from Baseline 2 and SpeechEmotionLlama v3 to an expressive speech prompt demonstrating this.
4.2 Speech emotion recognition
To more explicitly examine how well each system’s speech tokens convey emotion/style information to the LLM, we prompted the different model variants to perform SER by prepending the same prompt as used in the SER training task before the speech tokens. Then, we checked whether the response generated by the LLM matched the true emotion/style label. The rightmost column of Table 1 shows the resulting accuracies; the only exceptions are the cascade models, for which we list the performance of the standalone SER model. As expected, the two baselines performed quite poorly. Meanwhile, fine-tuning on emotion-conditioned response alignment resulted in a significant increase in performance, and adding speech captioning and SER each resulted in further improvements. Surprisingly, SpeechEmotionLlama v3 even slightly outperformed the standalone SER model, achieving 81.51% accuracy. This further demonstrates that our training framework allows the speech encoder to effectively convey paralinguistic information to the LLM in a way that allows the LLM to understand and make use of that information.
@@ 4.2 Speech emotion recognition
4.3 Analysis of speech tokens
We performed analyses to study the paralinguistic information encoded in the different systems’ speech tokens. We first trained linear probes on temporally mean-pooled speech tokens from each model variant to perform the 15-way SER task. The probes for the two baselines obtained accuracies of 25.08% and 31.20%, while those for SpeechEmotionLlama v1, v2, and v3 obtained higher accuracies of 68.15%, 73.15%, and 78.40%, respectively. We also performed a t-SNE [34] analysis of the mean-pooled speech tokens. Figure 2 displays the plots, which show that speech tokens from SpeechEmotionLlama are clustered much more clearly by emotion/style tag compared to the baseline. These results demonstrate that training on emotion-related response alignment tasks produces token embeddings that encode more discriminative paralinguistic information from speech.
5 Conclusion
We study the potential of an LLM to understand paralinguistic aspects of speech without any fine-tuning of its weights. Building on prior work that equips frozen LLMs with the ability to process speech inputs, we propose a novel training paradigm called behavioral alignment against emotion-conditioned responses. Our framework enables a speech encoder to generate token embeddings that capture both linguistic and paralinguistic information from spoken utterances and convey them to an LLM, even when the LLM remains completely frozen. The LLM, in turn, is able to effectively take the paralinguistic information in the speech tokens into account and provide higher quality and more empathetic responses to expressive speech prompts. Our work opens up possibilities for developing interfaces to LLMs that can leverage their capabilities for other kinds of speech or audio tasks without the need for fine-tuning, which can often be desirable for existing deployment settings.